# On the Fundamental Limitations of Dual Static CVaR Decompositions in Markov Decision Processes

### v1

@@ Abstract
Recent work has shown that dynamic programming (DP) methods for finding static CVaR-optimal policies in Markov Decision Processes (MDPs) can fail when based on the dual formulation, yet the root cause for the failure has remained unclear. We expand on these findings by shifting focus from policy optimization to the seemingly simpler task of policy evaluation. We show that evaluating the static CVaR of a given policy can be framed as two distinct minimization problems. For their solutions to match, a set of “risk-assignment consistency constraints” must be satisfied, and we demonstrate that the intersection of the constraints being empty is the source of previously observed evaluation errors. Quantifying the evaluation error as the CVaR evaluation gap, we then demonstrate that the issues observed when optimizing over the dual-based CVaR DP are explained by the returned policy having a non-zero CVaR evaluation gap. We then leverage our proposed risk-assignment perspective to prove that the search for a single, uniformly optimal policy via on the dual CVaR decomposition is fundamentally limited, identifying an MDP where no single policy can be optimal across all initial risk levels.
@@ 1 Introduction
The goal of reinforcement learning (RL) (Sutton and Barto, 2018) is to learn (sequential) decision-making policies such as to maximize some outcome (return) in a given environment, typically modeled as a Markov decision process (MDP). This is usually approached from the objective of maximizing the expected return, which has led to impressive successes in videogames (Vinyals et al., 2019), board games (Silver et al., 2018), and content recommendation (Li et al., 2010). However, in safety-critical domains like healthcare, autonomous driving, or financial planning, some erroneous actions may lead to disastrous consequences. For instance, in the task of identifying the shortest path to an organ for surgery, paths at high risk of endangering the patient (e.g., as they are too close to an artery, a nerve, or a critical region of the brain), should be avoided (Baek et al., 2018). Automation in this context and other safety-critical domains can only be achieved with safe decision-making policies (Gottesman et al., 2019). This can be achieved by optimizing a risk-averse objective instead of simply maximizing the expected return (Artzner et al., 1999). In particular, conditional value-at-risk (CVaR), which is considered a gold standard risk measure in banking regulations (Basel Committee on Banking Supervision, 2019), has received a lot of focus in risk-averse RL (Prashanth et al., 2022).
@@
More specifically, the static CVaR evaluation consists of computing the CVaR of a policy’s cumulative random return. Unfortunately, optimizing a policy w.r.t. the static CVaR objective in MDPs has proven to be quite challenging since CVaR suffers from the time inconsistency issue (Pflug and Pichler, 2016; Gagne and Dayan, 2021) and optimal policies may be history-dependent (Shapiro et al., 2014). To tackle both these problems, prior work has considered dynamic programs (DPs) applied on augmented state spaces in both the primal and dual representations of risk measures. Working under the primal representation, states are augmented by keeping track of the running cumulative return (Bäuerle and Ott, 2011; Boda et al., 2004; Chow and Ghavamzadeh, 2014; Chow et al., 2018). However, these methods are considered practically inefficient since they require computing the value function on an unbounded continuous state space (Chapman et al., 2021; Chow et al., 2015; Li et al., 2022). The dual representation has therefore been identified as a promising direction, where sequential decomposition of risk measures can also be leveraged (Chow et al., 2015; Pflug and Pichler, 2016). States here are augmented by keeping track of the current risk-level (between 0 and 1). Although the resulting augmented state space is still continuous, discretization can be applied efficiently since risk-levels are bounded (Chow et al., 2015; Li et al., 2022). Hence, Chow et al. (2015) proposed a Value Iteration (VI) procedure for CVaR in the dual representation, which served as the basis for many later developments in the field (Chapman et al., 2019, 2021; Chow et al., 2015; Ding and Feinberg, 2022a, b; Stanko and Macek, 2019; Rigter et al., 2021), until Hau et al. (2023) later showed counterexample MDPs where this procedure fails to recover the optimal value function and policy. However, from these few empirical results alone, it is not possible to understand what makes CVaR VI fail, and therefore how, or even if, it can be fixed.
@@
The main goal of this work is to diagnose the root causes of recently discovered failures in the static dual CVaR DP decomposition. To this end, we first establish a formal analysis framework that recasts the static CVaR evaluation and its DP decomposition as two distinct optimization problems over perturbations. Leveraging this perspective, we identify a set of risk-assignment consistency constraints that must be satisfied for the DP’s evaluation of a policy to be accurate. We show that the previously observed cases where the DP decomposition returned a suboptimal policy are explained by these constraints being inconsistent for the returned policy. We then leverage this constraint-based view to present an MDP where the action constraints required for optimality at different initial risk levels are irreconcilable. This proves that no single risk-dependent policy can be uniformly optimal, revealing a fundamental limitation of the pursuit of a single, universally optimal policy for all risk levels via the dual decomposition, independent of the DP algorithm used. To increase readability, we postpone most of the proofs to the appendix.
@@ 2 Static Risk-Averse Reinforcement Learning
Following Puterman (2014), let us define a finite Markov decision process (MDP) by a tuple $(\mathcal{S},\mathcal{A},P,\mathcal{R},s_{0},\gamma)$ where $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite action space, $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ is the transition function between states ($\Delta$ denoting the probability simplex), $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,R_{\text{max}}]$ is the reward function, $s_{0}\in\mathcal{S}$ is the initial state from which the process begins, and $\gamma\in[0,1)$ is a discount factor. At each time step $t\in\mathbb{N}_{0}$, an agent performs an action $A_{t}\in\mathcal{A}$ in the current state $S_{t}\in\mathcal{S}$ according to some decision policy $\pi$. This leads to a transition into the state $S_{t+1}$ sampled from $P(S_{t},A_{t})$, following which the agent receives the reward $R_{t+1}=\mathcal{R}(S_{t},A_{t},S_{t+1})$. Throughout this paper, we will adopt the convention of using uppercase letters to distinguish objects subject to random realization, such as referring to the state $S_{t}=s\in\mathcal{S}$.
Remark 1
@@
We consider policies as deterministic action-selection mechanisms. The most general form of policies is history-dependent policies $\pi_{h}:\mathcal{H}\to A$, where $\mathcal{H}$ is the set of histories defined as all previous states and actions seen prior to the action selection being made. Let $H_{t}\coloneq(S_{0},A_{0},S_{1},A_{1},S_{2},\dots,S_{t})$ and $\mathcal{H}_{t}$ respectively denote the current history and the set of possible histories at time $t$. At time $t=0$, the set of histories is limited to the initial state, that is $\mathcal{H}_{0}\coloneq\{(s_{0})\}$. For subsequent steps it is defined recursively as the combination of possible previous histories with the previous action and current state concatenated, that is $\mathcal{H}_{t+1}\coloneq\mathcal{H}_{t}\times\mathcal{A}\times\mathcal{S}$. Allowing a slight abuse of notation, we extend transition and action selection dynamics to histories and define the probability of observing a given history $H_{t}$ given policy $\pi_{h}$ as
$P^{\pi_{h}}(H_{t})\coloneq\prod_{\tau=0}^{t-1}P(S_{\tau+1}|S_{\tau},A_{\tau})\mathds{1}{\left[{A_{\tau}=\pi_{h}(H_{\tau})}\right]},$
@@
Because of the degree of complexity brought upon by policies operating on the possibly immense set of histories, we are often interested in Markovian policies $\pi:\mathcal{S}\to\mathcal{A}$, a special case of history-dependent policies where actions are selected only based on the current state. Hereafter, $\pi$ will be used to denote Markovian policies while $\pi_{h}$ will consider history-dependent policies.
@@
The return associated with a history $H$ in the MDP is defined as the discounted sum of rewards
@@
where $\gamma\in[0,1)$ is the MDP’s discount factor and $R_{t+1}=\mathcal{R}(S_{t},A_{t},S_{t+1})$. Given that trajectories in an MDP are generated using a random process, due to next state transitions being stochastic, we denote the random return of a trajectory generated by policy $\pi_{h}$ as a random variable $Z^{\pi_{h}}$, taking value $\mathcal{R}^{H}_{0:T}$ (Eq. 1) with $H\sim P^{\pi_{h}}$. The standard RL objective (Sutton and Barto, 2018) is to identify the optimal policy $\pi^{\star}_{h}$ that maximizes the expected return over histories
@@
It is known that Equation 2 can always be solved by a Markovian policy $\pi$ (Szepesvári, 2022). Unfortunately, because the expectation only weighs random outcomes according to their likelihood without taking their value into account, the optimal policy according to this objective may lead to catastrophic outcomes over some trajectories (Mannor et al., 2007). In critical applications where such trajectories should be avoided, one may optimize a risk-averse objective instead, where large negative outcomes are assigned higher importance.
@@ 2.1 Static CVaR for risk aversion
Let $Z\in\mathbb{R}$ denote a bounded variable on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, with cumulative distribution function $F_{Z}(z)=\mathbb{P}[Z\leq z]$ for some threshold $z\in\mathbb{R}$. Denote the Value-at-Risk (VaR) at risk-level $\alpha\in[0,1]$ as $\text{VaR}_{\alpha}[Z]\coloneq\min\left\{z\,|\,F_{Z}(z)\geq\alpha\right\}$. Assuming that $Z$ represents a payoff that should be maximized, the conditional-value-at-risk (CVaR) (Rockafellar and Uryasev, 2000; Föllmer and Schied, 2016) at risk level $\alpha$ is given by
[…]
The dual formulation in Equation 3 shows that the CVaR can be expressed as an optimization problem, where the objective is to find perturbations $\xi$ applied to the stochastic generative process of variable $Z$ such as to minimize its expectation. From the definition of the risk envelope $\Xi_{\alpha}\left(\mathbb{P}\right)$, we can observe that the perturbations enjoy two interesting properties. First, because $\xi$ represents multiplicative interventions on an event’s likelihood, it only affects events with nonzero probability. Also, because the largest magnitude of the perturbation on an event is $\frac{1}{\alpha}$, one can view $\frac{1}{\alpha}$ as a perturbation budget which is naturally minimal at $\alpha=1$ and increases as $\alpha$ decreases to 00, simultaneously recovering the monotonically increasing property of CVaR and its edge cases.
@@
[…]
@@ 2.2 CVaR-RL dynamic decomposition
Trajectory-level computation of static CVaR (Eq. 4) is impractical because it requires computing $P^{\pi_{h}}$ for all trajectories, which can be prohibitive for large state and action spaces. Fortunately, the CVaR decomposition Theorem (Pflug and Pichler, 2016; Chow et al., 2015) grants a recipe for expressing the evaluation at state-level.
Theorem 2 (CVaR decomposition, Thm. 2 from Chow et al. (2015))
@@
where $\tilde{\xi}$ is a perturbation over next state transitions, action $A_{t}$ is given by policy $\pi_{h}(H_{t})$, and $H^{\prime}=H_{t}\cup(A_{t},s^{\prime})$ is a possible history realization at time $t+1$.
Remark 3
We distinguish perturbations over next states ($\tilde{\xi}$) from perturbations over histories ($\xi$). While both perturbations impact the sampling of events, $\tilde{\xi}$ also dictates the risk-level evolution. This distinction is crucial and will be a core component of our analysis.
@@
Theorem 2 shows that the $\text{CVaR}_{\alpha}$ at any given time $t$ can be expressed as combination of $\text{CVaR}_{\alpha^{\prime}}$ values of possible next states $s^{\prime}$ for updated risk-levels $\alpha^{\prime}=\alpha\cdot\xi(s^{\prime})$ at time $t+1$. The running risk-level therefore contains all the information necessary to compute the CVaR of a history-dependent policy $\pi_{h}$. This motivated Chow et al. (2015) to introduce the risk-augmented state space $\tilde{\mathcal{S}}:\mathcal{S}\times[0,1]$, defined for any state $s\in\mathcal{S}$ and risk-level $y\in[0,1]$, and the corresponding risk-dependent Markovian policies on the augmented state space $\tilde{\pi}:\mathcal{S}\times[0,1]\to\mathcal{A}$. Chow et al. (2015) suggested that operating over $\tilde{\mathcal{S}}$ would suffice to retrieve the optimal history-dependent policy and evaluate its corresponding static CVaR through a value function mimicking the mechanism of Theorem 2.
Definition 4 (Risk-dependent-policy value function)
@@
where $a=\tilde{\pi}(s,y)$ is the action selected by the policy, $y^{\prime}=y\cdot\tilde{\xi}(s^{\prime})$ is the subsequent risk-level, and $\bm{\upsilon}^{\tilde{\pi}}_{0}(s,y)=0$ for all states $s\in\mathcal{S}$ and risk-levels $y\in[0,1]$. We let $\bm{\upsilon}^{\tilde{\pi}}\coloneq\bm{\upsilon}^{\tilde{\pi}}_{T}$.
Crucially, any risk-dependent policy $\tilde{\pi}$ induces a corresponding history-dependent policy $\tilde{\pi}_{h}^{\alpha}$ for a given initial risk-level $Y_{0}=\alpha$. At time $t$, given the risk-augmented state $(S_{t},Y_{t})$ and the selected action $A_{t}=\tilde{\pi}(S_{t},Y_{t})$, the subsequent risk-level is updated to $Y_{t+1}=Y_{t}\cdot\tilde{\xi}^{\star}(S_{t+1}|S_{t},Y_{t},A_{t})$, where the optimal perturbation $\tilde{\xi}^{\star}$ denotes the solution to the value function $\bm{\upsilon}^{\tilde{\pi}}(S_{t},Y_{t})$. Repeating this process $t$ times allows one to compute the action $\tilde{\pi}_{h}^{\alpha}(H_{t})$ for any history $H_{t}$. In the remainder of this paper, we will slightly abuse terminology and directly refer to a risk-dependent policy’s static CVaR to represent the static CVaR of its history-dependent counterpart.
In light of this correspondence between risk-dependent and history-dependent policies, Chow et al. (2015) proposed a Value Iteration algorithm, which we will refer to as CVaR VI, to find the risk-dependent policy with the optimal value function $\tilde{\pi}^{\star}\in\arg\max_{\tilde{\pi}}\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)$. Tentative proofs (Chow et al., 2015; Li et al., 2022) claimed $\tilde{\pi}^{\star}$ represented a risk-dependent version of the CVaR-optimal history-dependent policy $\pi_{h}^{\star}$, hence presenting a dynamic program decomposition of the CVaR-RL objective (Eq. 5). The optimality of the policy returned by CVaR VI was however refuted by Hau et al. (2023), who presented a counterexample MDP where the algorithm returns a suboptimal policy.
@@
In this work, we aim to explain why CVaR VI fails at a more fundamental level. The root cause we identify is the discrepancy between the value function of a risk-dependent policy and its corresponding static CVaR, which we formally define as the CVaR evaluation gap

### v2

@@ Abstract
It was recently shown that dynamic programming (DP) methods for finding static CVaR-optimal policies in Markov Decision Processes (MDPs) can fail when based on the dual formulation, yet the root cause of this failure remains unclear. We expand on these findings by shifting focus from policy optimization to the seemingly simpler task of policy evaluation. We show that evaluating the static CVaR of a given policy can be framed as two distinct minimization problems. We introduce a set of “risk-assignment consistency constraints” that must be satisfied for their solutions to match and we demonstrate that an empty intersection of these constraints is the source of previously observed evaluation errors. Quantifying the evaluation error as the CVaR evaluation gap, we demonstrate that the issues observed when optimizing over the dual-based CVaR DP are explained by the returned policy having a non-zero CVaR evaluation gap. Finally, we leverage our proposed risk-assignment constraints perspective to prove that the search for a single, uniformly optimal policy on the dual CVaR decomposition is fundamentally limited, identifying an MDP where no single policy can be optimal across all initial risk levels.
@@ 1 Introduction
The goal of reinforcement learning (RL) (Sutton and Barto, 2018) is to learn (sequential) decision-making policies such as to maximize some outcome (return) in a given environment, typically modeled as a Markov decision process (MDP). This is usually approached from the objective of maximizing the expected return, which has led to impressive successes in games (Silver et al., 2018; Vinyals et al., 2019) and content recommendation (Li et al., 2010). However, in safety-critical domains like healthcare, autonomous driving, or financial planning, some erroneous actions may lead to disastrous consequences. For instance, in the task of identifying the shortest path to an organ for surgery, paths at high risk of endangering the patient (e.g., as they are too close to an artery, a nerve, or a critical region of the brain), should be avoided (Baek et al., 2018). Automation in this context and other safety-critical domains therefore requires safe decision-making policies (Gottesman et al., 2019). This can be achieved by optimizing a risk-averse objective instead of simply maximizing the expected return (Artzner et al., 1999). In particular, conditional value-at-risk (CVaR), which is considered a gold standard risk measure in banking regulations (Basel Committee on Banking Supervision, 2019), has received a lot of focus in risk-averse RL (Prashanth et al., 2022).
@@
More specifically, the static CVaR evaluation consists of computing the CVaR of a policy’s cumulative random return. Unfortunately, optimizing a policy w.r.t. the static CVaR objective in MDPs has proven to be quite challenging since CVaR suffers from time inconsistency (Pflug and Pichler, 2016; Gagne and Dayan, 2021) and optimal policies may be history-dependent (Shapiro et al., 2014). To tackle these problems, prior work has considered dynamic programs (DPs) applied on augmented state spaces in both the primal and dual representations of risk measures. Working under the primal representation, states are augmented by keeping track of the running cumulative return (Boda et al., 2004; Bäuerle and Ott, 2011; Chow and Ghavamzadeh, 2014; Chow et al., 2018). However, such primal-based methods are considered practically inefficient since they require computing the value function on an unbounded continuous state space (Chow et al., 2015; Chapman et al., 2021; Li et al., 2022). The dual representation has therefore been identified as a promising direction, where the sequential decomposition of risk measures can be leveraged (Chow et al., 2015; Pflug and Pichler, 2016). States here are augmented by keeping track of the current risk level (between 0 and 1). Although the resulting augmented state space is still continuous, discretization can be applied efficiently since risk levels are bounded (Chow et al., 2015; Li et al., 2022). Hence, Chow et al. (2015) proposed a Value Iteration (VI) procedure for CVaR in the dual representation, which served as the basis for many later developments in the field (Chow et al., 2015; Chapman et al., 2019; Stanko and Macek, 2019; Chapman et al., 2021; Rigter et al., 2021; Ding and Feinberg, 2022b; a), until Hau et al. (2023) showed counterexample MDPs where this procedure fails to recover the optimal policy. However, from these few empirical results alone, it is not possible to understand what makes CVaR VI fail, and therefore how, or even if, it can be fixed.
@@
The main goal of this work is to diagnose the root causes of recently observed failures in the static dual CVaR DP decomposition. To this end, we first establish a formal analysis framework that recasts the static CVaR evaluation and its DP decomposition as two distinct optimization problems over perturbations. Leveraging this perspective, we identify a set of risk-assignment consistency constraints that must be satisfied for the DP evaluation of a policy to be accurate. We show that the previously observed cases where the DP decomposition returned a suboptimal policy are explained by these constraints being inconsistent for the returned policy. We then leverage this constraint-based view to present an MDP where the action constraints required for optimality at different initial risk levels are irreconcilable. This proves that no single risk-dependent policy can be uniformly optimal, revealing a fundamental limitation of the pursuit of a single, universally optimal policy for all risk levels via the dual decomposition, independent of the DP algorithm used. Practically, these findings suggest that the standard approach of training a single universal policy on the risk-augmented state space is structurally flawed. Our results indicate that practitioners should instead favor training specific policies for targeted risk profiles, akin to primal-based decomposition methods. To increase readability, we postpone most of the proofs to the appendix.
@@ 2 Static Risk-Averse Reinforcement Learning
Following Puterman (2014), let us define a finite Markov decision process (MDP) by a tuple $(\mathcal{S},\mathcal{A},P,\mathcal{R},s_{0},\gamma)$ where $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite action space, $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ is the transition function between states ($\Delta$ denoting the probability simplex), $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,R_{\text{max}}]$ is the reward function, $s_{0}\in\mathcal{S}$ is the initial state from which the process begins, and $\gamma\in[0,1)$ is a discount factor. At each time step $t\in\mathbb{N}_{0}$ of a trajectory, an agent performs an action $A_{t}\in\mathcal{A}$ in the current state $S_{t}\in\mathcal{S}$ according to some decision policy $\pi$. This leads to a transition into the state $S_{t+1}$ sampled from $P(S_{t},A_{t})$, following which the agent receives the reward $R_{t+1}=\mathcal{R}(S_{t},A_{t},S_{t+1})$. This process is repeated over an horizon of $T$ time steps. Throughout this paper, we adopt the convention of using uppercase letters to distinguish objects subject to random realization, such as referring to the state $S_{t}=s\in\mathcal{S}$. A table of notations can be found in Appendix A.
Remark 1.
@@
Assumptions 1.
Throughout this work, we consider finite state and action spaces ($|\mathcal{S}|<\infty,|\mathcal{A}|<\infty$). We focus on the finite-horizon setting with horizon $T$ and we include a discount factor $\gamma\in[0,1)$ to remain consistent with general formulations. Crucially, we restrict our attention to deterministic policies. This restriction is standard in the dual CVaR decomposition literature (Chow et al., 2015; Hau et al., 2023) to isolate external risk (stochasticity of the environment) from internal risk (policy randomization), simplifying the analysis of the risk level updates.
@@
We consider agent policies as deterministic action-selection mechanisms. The most general form of policies is history-dependent policies $\pi_{h}:\mathcal{H}\to A$, where $\mathcal{H}$ is the set of histories defined as all previous states and actions encountered prior to the current action-selection. Let $H_{t}\coloneq(S_{0},A_{0},S_{1},A_{1},S_{2},\dots,S_{t})$ and $\mathcal{H}_{t}$ respectively denote the current history and the set of possible histories at time $t$. At time $t=0$, the set of histories is limited to the initial state, that is $\mathcal{H}_{0}\coloneq\{(s_{0})\}$. For subsequent steps it is defined recursively as the combination of possible previous histories with the previous action and current state concatenated, that is $\mathcal{H}_{t+1}\coloneq\mathcal{H}_{t}\times\mathcal{A}\times\mathcal{S}$. Allowing a slight abuse of notation, we extend transition and action-selection dynamics to histories and define the probability of observing a given history $H_{t}$ given policy $\pi_{h}$ as
$P^{\pi_{h}}(H_{t})\coloneq\prod_{\tau=0}^{t-1}P(S_{\tau+1}|S_{\tau},A_{\tau})\mathds{1}{\left[{A_{\tau}=\pi_{h}(H_{\tau})}\right]},$ (1)
@@
Because of the degree of complexity brought upon by policies operating on the possibly immense set of histories, we are often interested in Markovian policies $\pi:\mathcal{S}\to\mathcal{A}$, a special case of history-dependent policies where actions are selected only based on the current state. Hereafter, $\pi$ will be used to denote Markovian policies while $\pi_{h}$ will denote history-dependent policies.
@@
The return associated with a history $H$ is defined as the discounted sum of rewards
@@
where $R_{t+1}=\mathcal{R}(S_{t},A_{t},S_{t+1})$. Given that trajectories in an MDP are generated using a random process, due to state transitions being stochastic, we denote the random return of a trajectory generated by policy $\pi_{h}$ as a random variable $Z^{\pi_{h}}$, taking value $\mathcal{R}^{H}_{0:T}$ (Eq. 2) with $H\sim P^{\pi_{h}}$ (Eq. 1). The standard RL objective (Sutton and Barto, 2018) is to identify the optimal policy $\pi^{\star}_{h}$ that maximizes the expected return over histories
@@
It is known that Equation 3 can always be solved by a Markovian policy $\pi$ (Szepesvári, 2022). Unfortunately, because the expectation only weights random outcomes according to their likelihood without taking their value into account, the optimal policy according to this objective may lead to catastrophic outcomes over some trajectories (Mannor et al., 2007). In critical applications where such trajectories should be avoided, one may optimize a risk-averse objective instead, where large negative outcomes are assigned higher importance.
@@ 2.1 Static CVaR for risk aversion
Let $Z\in\mathbb{R}$ denote a bounded variable on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, with cumulative distribution function $F_{Z}(z)=\mathbb{P}[Z\leq z]$ for some threshold $z\in\mathbb{R}$. Denote the Value-at-Risk (VaR) at risk level $\alpha\in(0,1]$ as $\text{VaR}_{\alpha}[Z]\coloneq\min\left\{z\,|\,F_{Z}(z)\geq\alpha\right\}$. Assuming that $Z$ represents a payoff that should be maximized, the conditional-value-at-risk (CVaR) (Rockafellar and Uryasev, 2000; Föllmer and Schied, 2016) at risk level $\alpha$ is given by
$\operatorname{CVaR}_{\alpha}\left[Z\right]\coloneq\frac{1}{\alpha}\int_{0}^{\alpha}\text{VaR}_{\beta}(Z)\text{d}\beta=\underbrace{\inf_{\xi\in\Xi_{\alpha}\left(\mathbb{P}\right)}\mathbb{E}_{\xi}\left[Z\right]}_{\text{dual formulation}},$ (4)
where $\Xi_{\alpha}\left(\mathbb{P}\right)\coloneq\left\{\xi:\omega\mapsto\left[0,\frac{1}{\alpha}\right]\,\Big|\,\int_{\omega\in\Omega}\xi(\omega)\mathbb{P}(\omega)d\omega=1\right\}$ defines the CVaRα risk envelope around distribution $\mathbb{P}$ and $\mathbb{E}_{\xi}[Z]$ is the $\xi$-reweighed expectation of $Z$. If $Z$ has a continuous distribution, it is well known that we have $\text{CVaR}_{\alpha}[Z]=\mathbb{E}\left[Z|Z\leq\text{VaR}_{\alpha}[Z]\right]$, which can be interpreted as the expected value of the worst $\alpha$ outcomes of $Z$. Note that CVaR is monotonically increasing in $\alpha$ with edge cases representing $\operatorname{CVaR}_{0}\left[Z\right]=\operatorname{ess}\inf[Z]$ and $\operatorname{CVaR}_{1}\left[Z\right]=\mathbb{E}[Z]$.
The dual formulation in Equation 4 shows that the CVaR can be expressed as an optimization problem, where the objective is to find perturbations $\xi$ applied to the stochastic generative process of variable $Z$ such as to minimize its expectation. From the definition of the risk envelope $\Xi_{\alpha}\left(\mathbb{P}\right)$, we can observe that the perturbations enjoy two interesting properties. First, because $\xi$ represents multiplicative interventions on an event’s likelihood, it only affects events with nonzero probability. Also, because the largest magnitude of perturbations on an event is $\frac{1}{\alpha}$, one can view $\frac{1}{\alpha}$ as a perturbation budget which is naturally minimal at $\alpha=1$ and increases as $\alpha$ decreases to 0, simultaneously recovering the monotonically increasing property of CVaR and its edge cases.
@@
where one aims to instead find a policy maximizing the CVaRα of its random return. Because CVaR[$_{\alpha}[Z^{\pi_{h}}$] can be intuitively interpreted as the expectation of the worst $\alpha$ fraction of trajectories when following policy $\pi_{h}$, optimizing the CVaR-RL objective should yield policies less prone to catastrophic outcomes than the standard RL objective (Eq. 3), with a lower $\alpha$ leading to increased cautiousness.
@@ 2.2 CVaR-RL dynamic decomposition
Trajectory-level computation of static CVaR (Eq. 5) is impractical because it requires computing $P^{\pi_{h}}$ for all trajectories, which can be prohibitive for large state and action spaces. Fortunately, the CVaR decomposition Theorem (Chow et al., 2015; Pflug and Pichler, 2016) grants a recipe for expressing the evaluation at state-level.
Theorem 1 (CVaR decomposition, Thm. 2 from Chow et al. (2015)).
@@
where $\tilde{\xi}$ are perturbations over next state transitions, action $A_{t}$ is given by policy $\pi_{h}(H_{t})$, and $H^{\prime}=H_{t}\cup(A_{t},s^{\prime})$ is a possible history realization at time $t+1$.
Remark 2.
We distinguish perturbations over next states ($\tilde{\xi}$) from perturbations over histories ($\xi$). While both perturbations impact the sampling of events, $\tilde{\xi}$ also updates the ongoing risk-level as dictated by the CVaR decomposition theorem (Thm. 1). When accumulated over an entire history, state perturbations implicitly yield history-level perturbations, but the connection between the two perturbation levels is complex and will be a core component of our analysis.
@@
Theorem 1 shows that the $\text{CVaR}_{\alpha}$ at any given time $t$ can be expressed as combination of $\text{CVaR}_{\alpha^{\prime}}$ values of possible next states $s^{\prime}$ for updated risk levels $\alpha^{\prime}=\alpha\cdot\tilde{\xi}(s^{\prime})$ at time $t+1$. The running risk level therefore contains all the information necessary to compute the CVaR of a history-dependent policy $\pi_{h}$. This motivated Chow et al. (2015) to introduce the risk-augmented state space $\tilde{\mathcal{S}}:\mathcal{S}\times(0,1]$, defined for any state $s\in\mathcal{S}$ and risk level $y\in(0,1]$, and the corresponding risk-dependent Markovian policies on the augmented state space $\tilde{\pi}:\tilde{\mathcal{S}}\to\mathcal{A}$. Chow et al. (2015) suggested that operating over $\tilde{\mathcal{S}}$ would suffice to retrieve the optimal history-dependent policy and evaluate its corresponding static CVaR through a value function based on the mechanism of Theorem 1.
Definition 1 (Risk-dependent-policy value function).
@@
where $a=\tilde{\pi}(s,y)$ is the action selected by the policy, $y^{\prime}=y\cdot\tilde{\xi}(s^{\prime})$ is the subsequent risk level, and $\bm{\upsilon}^{\tilde{\pi}}_{0}(s,y)=0$ for all states $s\in\mathcal{S}$ and risk levels $y\in(0,1]$. We let $\bm{\upsilon}^{\tilde{\pi}}\coloneq\bm{\upsilon}^{\tilde{\pi}}_{T}$.
Crucially, any risk-dependent policy $\tilde{\pi}$ induces a corresponding history-dependent policy $\tilde{\pi}_{h}^{\alpha}$ for a given initial risk level $Y_{0}=\alpha$. At time $t$, given the risk-augmented state $(S_{t},Y_{t})$ and the selected action $A_{t}=\tilde{\pi}(S_{t},Y_{t})$, the subsequent risk level is updated to $Y_{t+1}=Y_{t}\cdot\tilde{\xi}^{\star}(S_{t+1}|S_{t},Y_{t},A_{t})$, where the optimal perturbations $\tilde{\xi}^{\star}$ denote the solution to the value function $\bm{\upsilon}^{\tilde{\pi}}(S_{t},Y_{t})$. Repeating this process $t$ times allows to compute the action $\tilde{\pi}_{h}^{\alpha}(H_{t})$ for any history $H_{t}$. In the remainder of this paper, we slightly abuse terminology and refer to a risk-dependent policy’s static CVaR to represent the static CVaR of its history-dependent counterpart.
In light of this correspondence between risk-dependent and history-dependent policies, Chow et al. (2015) proposed a Value Iteration algorithm, which we refer to as CVaR VI, to find the risk-dependent policy with the optimal value function $\tilde{\pi}^{\star}\in\arg\max_{\tilde{\pi}}\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)$. Tentative proofs (Chow et al., 2015; Li et al., 2022) claimed $\tilde{\pi}^{\star}$ represented a risk-dependent version of the CVaR-optimal history-dependent policy $\pi_{h}^{\star}$, hence presenting a dynamic program decomposition of the CVaR-RL objective (Eq. 6). The optimality of the policy returned by CVaR VI was however refuted by Hau et al. (2023), who presented a counterexample MDP where the algorithm returns a suboptimal policy.
@@
In this work, we aim to explain why CVaR VI fails at a more fundamental level. To this end, we focus on the root cause, that is the discrepancy between the value function of a risk-dependent policy and its corresponding static CVaR, which we formally define as the CVaR evaluation gap